if angle A=90 2[r+R]=?

if angle A=90
2[r+R]=?

Grade:9

2 Answers

Arun Kumar IIT Delhi
9 years ago
Hello Student,

\\R=a/2 \\r=(s-c)tan({A \over 2})=s-c \\=>2(r+R)=2(s-c+c/2) \\=>b+c
Thanks & Regards
Arun Kumar
Btech, IIT Delhi
Askiitians Faculty
Leo Pixie
9 years ago
Amazing sir :)
Thanks